A charge Q on a conductor at time t is given by:-
Q = Q°(1 - e^-t/T)
Where Q° and T are constants.
a) Find the current I at any time t and
b) Sketch the variations of I and Q with time.
1
Expert's answer
2019-12-27T08:55:18-0500

For current

"I=\frac{dq}{dt}=\frac{Q_0 }{T}e^{-\frac{t}{T}}"